OpenStudy (australopithecus):

find a power series representation for the following function d/dx ln(5-x) = -1/5(1-(x/5)) so, \[-\int\limits_{}^{}\frac{dx}{5(1-\frac{x}{5})} = \sum_{n=0}^{\infty} \int\limits_{}^{}(\frac{x}{5})^{n}dx\] = \[\ln(5-x) + c = \sum_{n=0}^{\infty} \frac{5^{n}x^{-n+1}}{n(1-n)}\] by setting x = 0, c = -ln(5) so, \[\ln(5-x) = \ln(5) + \sum_{n=0}^{\infty} \frac{5^{n}x^{-n+1}}{n(1-n)}\] what did I do wrong it says this is incorrect. =

OpenStudy (australopithecus):

sorry it should be \[\ln(5-x) = \ln(5)- \sum_{n=0}^{\infty} \frac{5^{n}x^{-n+1}}{n(1-n)}\] I missed the negative

OpenStudy (australopithecus):

ugh I took the integral wrong its way too hot out to be doing math ugh

OpenStudy (australopithecus):

\[\ln(5-x) =\ln(5) - \sum_{n=0}^{\infty} \frac{x^{n+1}}{5^{n+1}(n+1)}\]

OpenStudy (australopithecus):

there it should be fixed now but it is still probably wrong

OpenStudy (anonymous):